OpenStudy (anonymous):

use elimination. -6x+5y=1 6x+4y=-10 is the answer (-1, -1)?

OpenStudy (amistre64):

dunno, can you show your steps to get there?

OpenStudy (amistre64):

it looks good

OpenStudy (anonymous):

correct!!!!

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

Check: -6(-1) + 5(-1) = 1 6(-1) + 4(-1) = -10 Yes, so it is correct!

OpenStudy (anonymous):

the 6x cancel each other. 5y+4y = 9y. 1 - 10 = -9. so you end up with 9y=-9 divide 9 and you get y= -1. substitute -1 in the first equation. -6x + 5(-1) =20. multiply 5*-1. now you have -6x - 5 = 1. add 5 to 1. -6x = 6. divide 6. 6/-6 = -1. therefor, my answer is (-1,-1) WHEW!

OpenStudy (amistre64):

very very good :)

OpenStudy (anonymous):

yay :) thanks